If h(x)=x^2+3x-1, find h(0)
Answer by jim_thompson5910(35256)



h(x)=x^2+3x-1

h(0)=(0)^2+3(0)-1

h(0)=0+3(0)-1

h(0)=0+0-1

h(0)=-1